OpenStudy (anonymous):

A fruit wingspantail tree grows oranges, grapefruits, and lemons. This season, there are 3 oranges, 6 grapefruits, and a number of lemons on the tree. If the probability of picking a lemon is 2/3, then how many lemons are on the fruit wingspantail tree? 18 that be 2/3 of 27

OpenStudy (shubhamsrg):

let there be x lemons total fruits = 3+6+x = 9+x probability for lemon = 2/3 = x/(9+x) solve for x

OpenStudy (anonymous):

i gotta get out more... what' a wingspantail tree?

OpenStudy (lgbasallote):

havent you seen one dp?

OpenStudy (anonymous):

I think i did actually... it was right next to Jacob's Ladder...

OpenStudy (lgbasallote):

yup that's the one

OpenStudy (anonymous):

i guess ppl focus on vocabulary as well thx

OpenStudy (lgbasallote):

he's just a troll forgive him...ive tried so many times to contain him

OpenStudy (anonymous):

sorry... :(

OpenStudy (anonymous):

whatever it is i still dont know whats a wingspanblabla

OpenStudy (anonymous):

the ratio that @shubhamsrg put up looks good.. just solve for x...